MONT 106Q – Mathematical Thinking Study/Discussion Questions on Othello October, 2016

Testing your comprehension and recall of the plot:

a. What is a “Moor”? What does the word mean? When does the action of the play take

place and what is happening in the affairs of state of Venice, Cyprus, and Turkey? b. Why does Desdemona love Othello as much as she does? Why does Othello feel worthy

of that love? c. What does Iago say about why he is so bent on harming Cassio and Othello? How does

he manipulate Roderigo, Cassio, and Othello himself to achieve his ends? How does he create doubt and jealousy in Othello? What literary devices such as symbolism, metaphors, etc. does Shakespeare have him use for this?

d. In Act I, how does Brabantio react to the news that Desdemona and Othello have married? How does Othello’s blackness (his race) feature in this? How is race discussed in the play with respect to Othello?

e. What is the role of Desdemona’s handkerchief in the action? What does it look like and how is it so easy to recognize? Where did she get it and what happens to it? Why is Othello so enraged by these happenings?

f. Compare and contrast Desdemona’s and Emilia’s thoughts about women and fidelity. What is Desdemona’s “Willow Song” and how does it foreshadow the ending of the play?

g. How does Othello approach the murder of Desdemona, and how does he justify it to himself? What symbols/images does Shakespeare employ? Why do you think Desdemona implies that she committed suicide? Why is she still protecting Othello? How does Emilia expose the truth?

h. Othello says he “loved not wisely but too well.” What does this mean? What happens to Othello and Iago at the very end?

Some larger themes:

a. In the ancient Greek tragic plays that defined the genre, a hero with the capacity for greatness but doomed by fate causes his own downfall through a tragic error (or perhaps a “flaw” – beware that this is a major point of disagreement between English literature specialists and Classics people who study the Greek plays(!)), while achieving some sort of moral victory at the end. Is Othello a tragedy in that sense? Are any of the defining characteristics from Greek drama true in Othello? If so, who is the tragic hero? What is his error or flaw?

b. Iago is probably the “meatiest” character in this play and famous male actors often chose to (or are selected to) play that part. What makes Iago such a challenging and attractive role to play? How are his speeches a mixture of verse and prose? Why do you think Shakespeare does that?

c. Why is it seemingly so easy for Iago to maneuver Othello from being a loving husband to being a jealous murderer (and in a remarkably short time too)? Is that quick change realistic or believable? Does social status and reputation have a role here? Are all the characters so dependent on their status and reputation for their happiness that they can be easily threatened? On the other hand, what makes Desdemona so steadfast in her love for her husband? Does the fact that she is a woman make a difference there?

d. Are Othello and Iago similar in some ways? If so, how? If not, why not? e. Are there parallels between this story and race relations in today’s society? f. What are some of the boundaries/borders that Shakespeare uses to structure his plot

and characters? Who has transcended those borders and how? Does transcending those borders end well or result in tragedy? What conclusions can we draw here?

Relating this with Othello in the Seraglio:

a. How does the opera Othello in the Seraglio draw on the Othello story? b. Does the Turkish musical setting add any new elements? What are they?